a rectangle has a width of 6 inches and a perimeter of 26 inches what is the length of the rectangle?
Answer by checkley77(12844)

2L+2W=26
2L+2*6=26
2L+12=26
2L=26-12
2L=14
L=14/2
L=7 ANS.
PROOF:
2*7+2*6=26
14+12=26
26=26